# Volatility Forecasting Using Similarity-based Parameter Correction and Aggregated Shock Information

### v2

@@ 3.2 Excess Volatility Estimators
The problem of aggregating estimated donor shocks begins with the data constraints. Taking the estimated shocks as a given, we essentially observe the pair ($(\{\hat{\omega}^{*}_{i}\}^{n+1}_{i=2},\{\textbf{v}_{i}\}^{n+1}_{i=2})$. We wish to recover weights $\{\pi_{i}\}^{n+1}_{i=2}\in\Delta^{n}$ leading to favorable forecasting properties. These weights are used to compute $\hat{\omega}^{*}\coloneq\sum^{n+1}_{i=2}\pi_{i}\hat{\omega}^{*}_{i}$, our forecast adjustment term. Since the weights $\{\pi_{i}\}_{i=2}^{n+1}$ are computed using $\mathcal{F}_{T^{*}_{i}}$, the set $\{\pi_{i}\}_{i=2}^{n+1}$ is deterministic, modulo any stochastic ingredient in the numerical methods employed to approximate $\textbf{v}_{1,T^{*}}$ using a convex combination of donor covariates. We say more about the properties of $\omega^{*}_{i}$ in section 4.
@@
In the forecast combination literature, it is discussed whether the weights employed to aggregate forecasts strive toward and meet various optimality criteria [Timmermann, 2006, Wang et al., 2023]. In this work, there are at least two senses of optimal weights that one might be interested in. First, we can think of optimal weights as a set $\{\pi_{i}\}_{i=2}^{n+1}$ such that $\omega_{1}=\sum^{n+1}_{i=2}\pi_{i}\hat{\omega}_{i}$, i.e., $\omega_{1}$ is recovered perfectly, as it belongs to convex hull of the estimated shocks. However, $\omega_{1}$ is never revealed to the practitioner, and hence there is no way of verifying the extent to which this condition is satifised.
@@ 4 Properties of Volatility Shocks and Shock Estimators
The model $\mathcal{M}_{1}$ is defined by a volatility equation and mean equation, as is any GARCH model. The choice to model the volatility shock $\omega^{*}_{i}$ as an additive random effect is straightforward. However, the choice to model the level effect $\epsilon^{*}_{i,t}$ as a temporary rupture in the otherwise i.i.d. sequence of innovations $\epsilon_{i,t}$ stands in need of deeper justification. One way of arguing for this choice is that, in a discrete time series model, if we assume the arrival of news in the time between $T^{*}$ and $T^{*}+1$, we do not have an easy way to express a conditional distribution of the innovation $\epsilon_{T^{*}+1}$ given the overnight arrival of information. Using $\epsilon^{*}_{i,t}$ thus breaks this impasse. This defense also explains why we do not parameterize the level shock at $T^{*}+1$ as a sum of two shocks, $\epsilon_{i,T^{*}+1}$ and $\epsilon^{*}_{i,T^{*}+1}$, which would represent the level shock as generated by two independent sources of stochasticity. To do so would be inelegant and would also lack motivation as a practical level. While we want to model the shock at $T^{*}+1$ as potentially large in absolute value, we also want to retain the property of a unitary source of noise.
Note that under the popular GARCH(1,1), a dual level-volatility shock has an marginal effect on the conditional variance $\sigma^{2}_{i,t}$ that reflects the geometric decay of innovations in autoregressive models. As usual, assume $\alpha+\beta<1$. Furthermore, assume that both the volatility shock $\omega^{*}_{i}$ and the level shock $\epsilon^{*}_{i,t}$ are of length one only, and consider a circumstance with no exogenous covariate $\textbf{v}_{i,t}$. Assume also that $r\geq 2$, which is necessary in order to isolate the effects of the level shock $\epsilon^{*}_{i,t}$. Then
@@
where $(\alpha(\mu^{2}_{\epsilon^{*}}+\sigma^{2}_{\epsilon^{*}})+\beta)$ may be greater than 1, permitting explosive behavior, at least in the short term. After both shocks have been exhausted, their influence disappears quickly. This short-memory effect has implications for the method being developed herein. First, there may be different risks associated with over/underestimating level shock and vol shock lengths. Estimation of effects in donor pool should err on the side of underestimating, not overestimating, the length of the max shock, since overestimation of the shock length brings with it the risk of underestimating $\omega^{*}$. Second, a practitioner of the method needs some idea of how long the operator expects the respective shocks in the time series under study to be. There are couple of obvious strategies: take all the donors, and over all the donor shock lengths, take the minimium. Alternatively, one could take the maximum.
@@ 4.1 Consistency of the fixed effect estimators
Under assumption 2, $\forall i,\{\omega_{i,t}^{*}\}_{t=0,...,T_{i}}$ is a strictly stationary series.
@@ 4.3 Asymptotic Loss
It is easily verified that as $\hat{\omega}^{*}_{t+1}\rightarrow 0^{+}$, the difference in the losses goes to zero. On the other hand, as $\hat{\omega}^{*}_{t+1}$ becomes large, the difference in the losses turns negative, with the lesson being that $\hat{\omega}^{*}_{t+1}$ must be in appropriate proportion to the volatility $\sigma^{2}_{t+1}$ in order for the adjusted forecast to outperform the unadjusted forecast. This explains why it is so important to avoid using a naive adjustment estimator, $\overline{\omega^{*}}$, the arithmetic mean of the estimated shocks. We conclude this section with a broader result.
@@ 5 Numerical Examples
In this section, we demonstrate the effectiveness of the proposed method using Monte Carlo simulations. In order to investigate the forecasting method presented herein, all of our simulations will use $\mathcal{M}_{1}$ volatility models and $\mathcal{M}_{0}$ models on the returns, i.e. $D^{return}_{i,t}\equiv 0$. We first explain the parameters to be varied, the parameters that remain fixed, and the behavior we expect to observe. For our purposes, a parameter is defined to be fixed whenever it does not vary across any of the simulations performed, and conversely a parameter is varying whenever it varies across at least two simulations performed.
@@ 5.2 Varying Parameters
We vary exactly five parameters: $\mu_{V},\sigma_{V},\mu_{\delta},\mu_{\omega^{*}},\sigma_{u}$, each of which is critical to evaluating the method’s responsiveness to changing conditions in the shock distribution. $\mu_{V},\sigma_{V}$ govern the elements of the volatility profile. $\mu_{\delta}$ interacts with $V$ via the dot-product operation, of course. $\mu_{\omega^{*}}$ is a location parameter for the volatility shock, and $\sigma_{u}$ is the idiosyncratic noise of the volatility shock.
@@ 5.3 Evaluative Framework and Monte Carlo Results
Lastly, the third set of simulations takes a look at the role of the intercept of the conditional shock distribution, $\mu_{\omega^{*}}$, in particular, its interaction with $\sigma_{u}$. We can think of $\mu_{\omega^{*}}$ as being the sole component of v that has no variance, i.e. because it’s corresponding entry in the vector v is the scalar 1. For larger values of $\mu_{\omega^{*}}$, the outperformance of the adjusted forecast declines with increase noise $\sigma_{u}$, but this effect is absent for smaller values of $\mu_{\omega^{*}}$
@@ 6 Real Data Example
We now discuss the three subplots in Figure LABEL:fig:SVF_2016 in order from left to right. On the left, we see that distanced-based weighting places nearly equal weight on the 2004 and 2012 elections, with only neglible weight on the 2008 election, when financial market conditions were extreme across nearly every dimension. Assuming an approximately correct specification of the covariates, this is interpreted to mean that even of the 2016 US election had a general climate of risk and tension less extreme than 2008 and more similar to the 2004 and 2012 elections. In the middle plot, we notice that the fixed effect estimates for 2008 and 2012 are considerable, with only 2004 registering a near-zero fixed effect estimate. The fixed effect estimates quantify the amount of surprise the US election results delivered (strictly speaking, not only the presidential race but all November elections in the US with the ability to influence financial markets) under the assumption of a GARCH(1,1). As estimates gleaned from only one data point per time series, they are theoretically high in variance. On the right, we observe in black the $\hat{\sigma}^{2}$ yielded by the GARCH(1,1) for the time series under study. We also observe four colored points, each listen in the legend: three predictions and the ground truth. We include the prediction derived by adjusting the GARCH(1,1) prediction by the arithmetic mean of the fixed effect estimates. As is evident, our method comes reasonably close the ground truth. The prediction is not only directionally correct, i.e. we predict a volatility spike where there is one; the prediction far outperforms the unadjusted prediction. Remarkably, the arithmetic-mean based prediction here demonstrates the inherent risk in failing to weight each donor appropriately. The 2008 election receives far more weight than is called for, as simple averaging ignores the radically different conditons on the evening of those two events.
@@ 7.4 Parameter Stability
There is an important question about the stability of the GARCH parameters under the presence of a shock, and on parameter instability as it pertains to forecasting, we refer readers to Rossi [2013, 2021]. There are at least two reasons that we do not herein explore parameter instability or methods to adjust for it. First, the marginal effect of coefficient changes at the shock time would, under the assumptions in this work, be swamped by the random effect. Second, the estimation of post-shock parameter values would require at least several — better yet, dozens — of post-shock data points, whereas this work assumes access to zero post-shock data points. However, it is possible that similarity-based estimators for the GARCH coefficients could be produced, for example, by adapting the methods of Dendramis et al. [2020].

### v3

@@ 3.2 Excess Volatility Estimators
The problem of aggregating estimated donor shocks begins with the data constraints. Taking the estimated shocks as a given, we essentially observe the pair ($(\{\hat{\omega}^{*}_{i}\}^{n+1}_{i=2},\{\textbf{v}_{i}\}^{n+1}_{i=2})$. We wish to recover weights $\{\pi_{i}\}^{n+1}_{i=2}\in\Delta^{n}$ leading to favorable forecasting properties. These weights are used to compute $\hat{\omega}^{*}\coloneq\sum^{n+1}_{i=2}\pi_{i}\hat{\omega}^{*}_{i}$, our forecast adjustment term. Since the weights $\{\pi_{i}\}_{i=2}^{n+1}$ are computed using $\mathcal{F}_{T^{*}_{i}}$, the set $\{\pi_{i}\}_{i=2}^{n+1}$ is deterministic, modulo any stochastic ingredient in the numerical methods employed to approximate $\textbf{v}_{1,T^{*}}$ using a convex combination of donor covariates. We say more about the properties of the shocks $\omega^{*}_{i}$ in section 4.
@@
In the forecast combination literature, it is of interest whether the weights employed to aggregate forecasts strive toward and meet various optimality criteria [Timmermann, 2006, Wang et al., 2023]. In our endeavor, there are at least two senses of optimal weights that one might be interested in. First, we can think of optimal weights as a set $\{\pi_{i}\}_{i=2}^{n+1}$ such that $\omega_{1}=\sum^{n+1}_{i=2}\pi_{i}\hat{\omega}_{i}$, i.e., $\omega_{1}$ is recovered perfectly, as it belongs to convex hull of the estimated shocks. However, $\omega_{1}$ is never revealed to the practitioner, and hence there is no way of verifying the extent to which this condition is satifised.
@@ 4 Properties of Volatility Shocks and Shock Estimators
The model $\mathcal{M}_{1}$ is defined by a volatility equation and mean equation, as is any GARCH model. The decision to model the volatility shock $\omega^{*}_{i}$ as an additive random effect is straightforward. However, the decision to model the level effect $\epsilon^{*}_{i,t}$ as a temporary rupture in the otherwise i.i.d. sequence of innovations $\epsilon_{i,t}$ stands in need of deeper justification. One way of arguing for this choice is that, in a discrete time series model, if we assume the arrival of news in the time between $T^{*}$ and $T^{*}+1$, we do not have an easy way to express a conditional distribution of the innovation $\epsilon_{T^{*}+1}$ given the overnight arrival of information. Using $\epsilon^{*}_{i,t}$ thus breaks this impasse. This justification also explains why we do not parameterize the level shock at $T^{*}+1$ as a sum of two shocks, $\epsilon_{i,T^{*}+1}$ and $\epsilon^{*}_{i,T^{*}+1}$, which would represent the level shock as generated by two independent sources of stochasticity. To do so would be inelegant and would also lack motivation as a practical level. While we want to model the shock at $T^{*}+1$ as potentially large in absolute value, we also want to retain the property of a unitary source of noise.
Note that under the popular GARCH(1,1), a dual level-volatility shock has an marginal effect on the conditional variance $\sigma^{2}_{i,t}$ that reflects the geometric decay of innovations in autoregressive models. As usual, assume $\alpha+\beta<1$. Furthermore, assume that both the volatility shock and the level shock are of length one only, and consider a circumstance with no exogenous covariate $\textbf{v}_{i,t}$. Assume also that $r\geq 2$, which is necessary in order to isolate the effects of the level shock $\epsilon^{*}_{i,t}$. Then
@@
where $(\alpha(\mu^{2}_{\epsilon^{*}}+\sigma^{2}_{\epsilon^{*}})+\beta)$ may be greater than 1, permitting explosive behavior, at least in the short term. After both shocks have been exhausted, their influence disappears quickly. This short-memory effect has implications for the method being developed herein. First, there may be different risks associated with over/underestimating level shock and vol shock lengths. Estimation of effects in donor pool should err on the side of underestimating, not overestimating, the length of the max shock, since overestimation of the shock length brings with it the risk of underestimating $\omega^{*}$. Second, a practitioner of the method needs some idea of how long the operator expects the respective shocks in the time series under study to last. There are couple of obvious strategies: take all the donors, and over all the donor shock lengths, take the minimium. Alternatively, one could take the maximum.
@@ 4.1 Consistency of the fixed effect estimators
Under assumption 2, for each $i,i=1,...,n+1$, $\{\omega_{i,t}^{*}\}_{t=0,...,T_{i}}$ is a strictly stationary series.
@@ 4.3 Asymptotic Loss
It is easily verified that as $\hat{\omega}^{*}_{t+1}\rightarrow 0^{+}$, the difference in the losses goes to zero. On the other hand, as $\hat{\omega}^{*}_{t+1}$ becomes large, the difference in the losses turns negative, with the lesson being that $\hat{\omega}^{*}_{t+1}$ must be in appropriate proportion to $\sigma^{2}_{t+1}$ in order for the adjusted forecast to outperform the unadjusted forecast. This explains why it is so important to avoid using a naive adjustment estimator, $\overline{\omega^{*}}$, the arithmetic mean of the estimated shocks. We conclude this section with a broader result.
@@ 5 Numerical Examples
In this section, we demonstrate the effectiveness of the proposed method using Monte Carlo simulations. All simulations will use $\mathcal{M}_{1}$ volatility models and $\mathcal{M}_{0}$ models on the returns, i.e. $D^{return}_{i,t}\equiv 0$ for each $i$ and each $t$. We first explain the parameters to be varied, the parameters that remain fixed, and the behavior we expect to observe. For our purposes, a parameter is defined to be fixed whenever it does not vary across any of the simulations performed, and conversely, a parameter is varying whenever it varies across at least two simulations performed.
The overarching story told by the simulations is that our method performs well as the magnitude of the signal in the shock grows in comparison to the idiosyncratic noise term of the shock. For brevity, we refer to this as the signal-to-noise phenomenon. However, this phenomenon is complicated somewhat by at least two nuances: additional parameters must be in place and properly sized in order to enable to signal-to-noise phenomenon, and second, certain hurdles must be met in the size of those additional parameters before the signal-to-noise story can be seen. In short, interaction effects abound, with non-linearities as well.
@@ 5.2 Varying Parameters
We vary exactly five parameters: $\mu_{V},\sigma_{V},\mu_{\delta},\mu_{\omega^{*}},\sigma_{u}$, each of which is critical to evaluating the method’s responsiveness to changing conditions in the shock distribution. $\mu_{V},\sigma_{V}$ govern the elements of the covariates. $\mu_{\delta}$ interacts with $\textbf{V}_{t}$ via the dot-product operation, of course. $\mu_{\omega^{*}}$ is a location parameter for the volatility shock, and $\sigma_{u}$ is the idiosyncratic noise of the volatility shock.
@@ 5.3 Evaluative Framework and Monte Carlo Results
Lastly, the third set of simulations takes a look at the role of the intercept of the conditional shock distribution, $\mu_{\omega^{*}}$, in particular, its interaction with $\sigma_{u}$. We can think of $\mu_{\omega^{*}}$ as being the sole component of v that has no variance, i.e. because it’s corresponding entry in the vector v is the scalar 1. For larger values of $\mu_{\omega^{*}}$, the outperformance of the adjusted forecast declines with increases in $\sigma_{u}$, but this effect is absent for smaller values of $\mu_{\omega^{*}}$
@@ 6 Real Data Example
Real Data Example: 2016 Election
Figure 5: The volatility induced by the 2016 US election
We now discuss the three subplots in Figure 5 in order from left to right. On the left, we see that distanced-based weighting places nearly equal weight on the 2004 and 2012 elections, with only neglible weight on the 2008 election, when financial market conditions were extreme across nearly every dimension. Assuming an approximately correct specification of the covariates, this is interpreted to mean that even of the 2016 US election had a general climate of risk and tension less extreme than 2008 and more similar to the 2004 and 2012 elections. In the middle plot, we notice that the fixed effect estimates for 2008 and 2012 are considerable, with only 2004 registering a near-zero fixed effect estimate. The fixed effect estimates quantify the amount of surprise the US election results delivered (strictly speaking, not only the presidential race but all November elections in the US with the ability to influence financial markets) under the assumption of a GARCH(1,1). As estimates gleaned from only one data point per time series, they are theoretically high in variance. On the right, we observe in black the $\hat{\sigma}^{2}$ yielded by the GARCH(1,1) for the time series under study. We also observe four colored points, each listen in the legend: three predictions and the ground truth. We include the prediction derived by adjusting the GARCH(1,1) prediction by the arithmetic mean of the fixed effect estimates. As is evident, our method comes reasonably close the ground truth. The prediction is not only directionally correct, i.e. we predict a volatility spike where there is one; the prediction far outperforms the unadjusted prediction. Remarkably, the arithmetic-mean based prediction here demonstrates the inherent risk in failing to weight each donor appropriately. The 2008 election receives far more weight than is called for, as simple averaging ignores the radically different conditons on the evening of those two events.
@@ 7.4 Parameter Stability
There is an important question about the stability of the GARCH parameters under the presence of a shock, and on parameter instability as it pertains to forecasting, we refer readers to Rossi [2013, 2021]. There are at least two reasons that we do not explore parameter instability or methods to adjust for it. First, the marginal effect of coefficient changes at the shock time would, under the assumptions in this work, be swamped by the random effect. Second, the estimation of post-shock parameter values would require at least several — better yet, dozens — of post-shock data points, whereas this work assumes access to zero post-shock data points. However, it is possible that similarity-based estimators for the GARCH coefficients could be produced, for example, by adapting the methods of Dendramis et al. [2020].